# Giant and robust Josephson diode effect in multiband topological nanowires

### v2

@@ Abstract
We also note that the balance is resulted from the competition between the Josephson current originating from the higher-curvature red band and that originating from the lower-curvature blue subbands. In general, through the spin-parity band exchange, we can combine the $N_{y}$ lower-curvature bands of identical spin parity to balance the $N_{y}-1$ opposite spin-parity bands with higher curvature. Such a mechanism provides a generic practical route to optimize the Josephson diode effect with high and robust efficiency in the multiband topological regime.
@@
Conclusion.–We have shown that multiband Majorana nanowires provide a realistic and robust platform for realizing the giant Josephson diode effect. The coexistence of MBS and ABS in the multiband regime leads to a persistent competition between the fractional and conventional Josephson currents, sustaining large diode efficiency deep into the topological phase. Especially, we uncovered a novel spin-parity band exchange mechanism unique to the multiband regime, which realizes a robust high efficiency plateau. These results demonstrate subband engineering as a practical route to optimize superconducting diode effect and offer a new probe for identifying topological phase in Majorana nanowire systems.
@@
Acknowledgments.– The authors are thankful to Yue Mao and Yu-Hao Wan for fruitful discussions. This work was supported by National Key Research and Development Program of China (2021YFA1400900), the National Natural Science Foundation of China (Grants No. 12425401, No. 12261160368, and No. 11921005), the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302000), and Shanghai Municipal Science and Technology Major Project (No. 2019SHZDZX01).
@@
For the possible values of $h_{x}$ and $h_{y}$, we note that the effective Land¨¦ factor in an InSb-based semiconducting nanowire is $g_{\text{eff}}\approx 50$, with $t_{x}$ and $\alpha_{x}$ comparable to those in InAs. By adopting a dimensionless $h$ and normalizing $\Delta$ to unity in the computation, the corresponding magnetic field in SI units is given by $B=\Delta/(g_{\text{eff}}\mu_{\text{B}})\cdot h=0.07h\cdot\text{T}$. Thus, achieving the high-efficiency regime requires a magnetic field just below $1\text{T}$, which is experimentally available.

### v3

@@ Abstract
We also note that this balance stems from the competition between the Josephson current originating from the higher-curvature red band (with lower chemical potential from band bottom) and that originating from the lower-curvature blue subbands (with higher chemical potential from band bottom). In general, through the spin-parity band exchange, we can combine the $N_{y}$ lower-curvature bands of identical spin parity to balance the $N_{y}-1$ opposite spin-parity bands with higher curvature. Such a mechanism provides a novel generic scheme to optimize the JDE with high and robust efficiency in the multiband topological regime.
@@
Conclusion.–We have shown that multiband Majorana nanowires provide a realistic and robust platform for realizing the giant JDE. The coexistence of MBS and ABS in the multiband regime leads to a persistent competition between the fractional and conventional Josephson currents, sustaining large diode efficiency deep into the topological phase. Especially, we uncovered a novel spin-parity band exchange mechanism unique to the multiband regime, which realizes a robust high efficiency plateau. These results demonstrate subband engineering as a practical route to optimize SDE and offer a new probe for identifying topological phase in Majorana nanowire systems.
@@
Acknowledgments.– The authors are thankful to Yue Mao and Yu-Hao Wan for fruitful discussions. This work was supported by the National Natural Science Foundation of China (Grants No. 12425401 and No. 12261160368), the National Key Research and Development Program of China (2021YFA1400900), Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2021ZD0302000), and Shanghai Municipal Science and Technology Major Project (No. 2019SHZDZX01).
@@
For the possible values of $h_{x}$ and $h_{y}$, we note that the effective Lande factor in an InSb-based semiconducting nanowire is $g_{\text{eff}}\approx 50$, with $t_{x}$ and $\alpha_{x}$ comparable to those in InAs. By adopting a dimensionless $h$ and normalizing $\Delta$ to unity in the computation, the corresponding magnetic field in SI units is given by $B=\Delta/(g_{\text{eff}}\mu_{\text{B}})\cdot h=0.07h\cdot\text{T}$. Thus, achieving the high-efficiency regime requires a magnetic field just below $1\text{T}$, which is experimentally available.